OpenStudy (anonymous):

Travis earned $42.50 in 5 h. How much will he earn in 8 h at the same rate? $?

OpenStudy (anonymous):

what do you think it is?

OpenStudy (texaschic101):

there are a couple of ways to do this...I will show you both (1) Divide 42.50 by 5 and that is what he makes per hr (2) Then multiply that answer by 8, and that will tell you how much he makes for 8 hrs. another way is to set it up as a proportion 42.50/5 = x/8 -- now cross multiply and then solve for x Both of these ways should get you the same answer

OpenStudy (anonymous):

42.5 / 5 =?

OpenStudy (anonymous):

answer is 68

OpenStudy (anonymous):

@Ihxza2 Thanks